OpenStudy (shelby1290):

For each quadratic relation, determine the y-intercept and the axis of symmetry a) y= x^2 + 2 b) y= -x^2 -4

OpenStudy (shelby1290):

For part A) I got y=2 because I set x to 0 to get my y-intercept y=x^2+2 y=0^2+2 y=2 (0,2)

OpenStudy (shelby1290):

I then set my y-int to 0 to get my x y=x^2+2 0=x^2+2 0=1x+2 x=2 (2,0)

OpenStudy (shelby1290):

Am I doing it right so far? How do i find the axis of symmetry?

OpenStudy (anonymous):